In a study of a common insect, the number of bristles on the fifth abdominal sternite was shown to follow a normal distributiI just need help with B.
Answer #1

Only answer for part (b) has been asked.

(b)We need an interval centered around the mean which covers 90% of male population. Let this interval be (L, U). The area of the distribution outside this interval has 100%-90%=10% of the population. SInce normal distriburtion is symmetric, this means that the part of the distribution from  \small -\infty to L has 5% of the population and part of the distribution from U to \small \infty has 5% of the population. Area between L and U has 90% of the population. Therefore area between \small -\infty and U has 5% + 90% = 95% of the population.

Let CDF be the cumulative density function of the Standard Normal distribution. Converting L and U to standard normal variables, we get Z1 = (L - 17.8)/1.7, Z2 = (U -17.8)/1.7

Hence CDF(Z1) = 0.05 and CDF(Z2) = 0.95.

Hence, Z1 = -1.645 and Z2 = 1.645

Or, (L - 17.8)/1.7 = -1.645 and (U - 17.8)/1.7 = 1.645

Or L = 15.0035 and U = 20.5965

Hence, the interval is 15.0 to 20.6 abdominal bristles.